# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 4.3 Semi-supervised approach
In this section we describe our specific framework for RESET’s Step 3, which as mentioned, can use any semi-supervised machine learning model. It consists of two main steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
  1. i.
    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo-targets (which we elaborate on in Step III at the end of this section). We define the features that are subsequently used by a collection of classification algorithms as the combined score and side information $(W,\mathbf{x})$.
  2. ii.
    For each considered classification algorithm, we randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout). For each $k\in\{1,2,\dots,K\}$, we train the classification algorithm using the positive and negative set on the features in the $K\setminus\{k\}$ folds and apply the trained model to all the hypotheses in the $k$th test fold. We consider a random forest (RF), a generalized additive model (GAM), and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
  3. iii.
    To reduce variability, we repeat Step ii $r$ times (we used $r=10$).
  4. iv.
    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1, which is the similar to SSS+ of Equation (2) without the ‘+1’) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels. If $\mathbb{P}(L_{i}=1)\leq c_{0}$ then we set $c=1-s\cdot(1-c_{0})$. This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo-targets to training decoys [^4] This is not a requirement for error control..
  5. v.
@@
  6. vi.
    Using the chosen classification algorithm, we assign new scores $\tilde{W}$ to the hypotheses. $\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis in the test fold that it appears in, taken over the $r$ repetitions executed in Step iii.
  7. vii.
    We reapply Selective SeqStep (SSS) at $\alpha$ using the pseudo labels and the updated scores, $(\tilde{L},\tilde{W})$ with $c=1-s\cdot(1-c_{0})$.
  8. viii.
    The positive set is redefined as those pseudo-targets that are ‘discovered’ in Step vii. Steps ii-vi are then repeated with the new positive set. To ensure that the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set. We used min_positive = 50.
  9. ix.
    The final scores $\tilde{W}$ are then reported.
We apply the following additional heuristics to our implementation that occur prior to Step 1 and leave some minor details in Section S.4.
  1. I.
    RESET throws out the hypotheses with undefined labels, i.e., those hypotheses with $W_{i}=0$. This effect is negligible when using p-values. However, in variable selection, the LCD score in the knockoff filter forces many of the winning scores to be zero ($W_{i}=0$), leaving typically a few hypotheses to be considered by RESET and consequently losing some information. We try to recover some of this information in the following way. For each hypothesis with $W_{i}\neq 0$, we identify the k-nearest neighbors in terms of the side information (zero-scoring hypotheses included). Then we define an extra side information variable as the number of zero-scoring hypotheses among these identified neighbors (we used the 20-nearest neighbors).
  2. II.
    To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure. That is, we use a generalized additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time. We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit. Again, we try to make use of some of the zero-scoring hypotheses setting $W\cdot\tilde{L}=0$ in these cases. We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
  3. III.
    Not all the pseudo-targets need to be included in the initial positive set in Step i. After all, we expect many of the pseudo-targets to be true nulls. Hence, to improve the performance of RESET, we consider the following strategy. For each feature in $(W,\mathbf{x})$, reorder the hypotheses according to that feature, and apply SSS to the pseudo labels $\tilde{L}$ at the FDR level $\alpha$. Then select the feature that maximizes the number of pseudo discoveries. Then using the chosen feature, we define the positive set as those pseudo-targets discovered at $\alpha_{0}$, again using SSS with the pseudo labels $\tilde{L}$ (we used $\alpha_{0}=50\%$ throughout). Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed, of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board). To ensure that the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set.
We refer to the general RESET wrapper in combination with the above semi-supervised approach as RESET Ensemble. We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step ii is reduced to a random forest, generalized additive model, and a collection of neural networks as described, respectively.
@@ 4.4 Code availability
Our code is available at https://github.com/freejstone/stat_RESET_paper_code with an Apache-2.0 license. We intend to develop a package for RESET in R.
@@ 5 Experiments: competition setting
5.1 Numerical experiments
@@
In this section, we conduct a wide range of simulations in the variable selection problem based on those introduced by Ren and Candès [44]. We compare RESET Ensemble to Adaptive Knockoffs (AdaKO) and the generic Knockoff Filter (KO) (see Section 2). Further implementation details for AdaKO are given in Supplementary Section S.5. We provide a brief description of each simulation setup, referring to [44] for further details. Each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. The labels and scores are obtained according to Equation (1). We focus on FDR control in this section, and we study FDP control in Section 5.2.
5.1.1 Simulation 1: Linear model with one-dimensional side information
Our first simulation is based on Simulation 1 by Ren and Candès [44]. They used $n=1000$ independent observations from the joint data distribution of $(X,Y)$, where $X$ is drawn from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,1)$ is a linear model. The $\beta$’s are selected in the following way. Let $k$ denote the number of nonzero $\beta$’s. Then $k$ indices from $\{1,2,\dots,p\}$ are sampled without replacement with indices $i\in\{1,\dots,2k\}$ having probability proportional to $1/i^{2}$ and the indices $i\in\{2k+1,\dots,p\}$ having zero probability. The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d. uniformly. We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$. Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s, because smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$. We used 100 independent runs of the above, where each run consisted of drawing those $n$ observations. The original simulation in [44] sampled the $\beta$’s once and fixed them across the 100 runs. Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed that all the methods exhibited significant variability between the draws.
5.1.2 Simulation 2: Linear model with two-dimensional side information
In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section 5.1.1, with $p=1800$. Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$. Note that unlike the previous simulation, it is not relevant how the $k$ variable indices are selected from $\{1,2,\dots,p\}$. Each hypothesis is equipped with the following two-dimensional side information. For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d. uniformly. We used 20 independent runs of this larger simulation instead of 100, where in each run we draw those $n$ observations, $\beta$ coefficients, and side information variables.
5.1.3 Simulation 3: Logistic model with two-dimensional side information
Next, we consider Simulation 2 by Ren and Candès [44]. They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$, where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [44] for further details). Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $[-20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$. The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice. An image of the nonzero $\beta$’s is given in Figure 1. Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d. uniformly. We used 98 runs of the above setup (since in 2 runs, AdaKO EM failed with an error), where each run consisted of drawing those $n$ observations. The original simulation in [44] sampled the sign of the $\beta$’s once, fixing them across the 100 runs. Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found that the results of all the methods significantly varied with those draws.
Figure 1: Location of nonzero $\beta$’s according to the pair of values $(x,y)$ in the lattice. The same image is given in [44].
Figure 2: Scatterplot of each hypothesis according to the LCD score denoted as $W$ (x-axis) and side information variable (y-axis). Each hypothesis is colored according to whether it is a true (red) or false null hypothesis (blue). In (A), the side information variable is the index of the hypothesis in Simulation 1 (Section 5.1.1). In (B), we use one of the three side information variables as described in Simulation 4 (Section 5.1.4).
5.1.4 Simulation 4: Large $p$ and $n$ and 3-d side information
Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1. Each variable is equipped with the following three-dimensional side information. The first $k$ variables are randomly assigned a value drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$. The remaining $p-k$ variables are uniquely assigned to one of the remaining values from $\{1,2,\dots,p\}$ uniformly at random. We expect most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1, as depicted in Figure 2. This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative. We conducted 5 independent runs, where in each run we drew the $n$ observations, $\beta$ coefficients, and side information.
@@ 5.1.5 Results
We first compare RESET Ensemble with Adaptive Knockoff’s default method in R, AdaKO GAM, and the generic Knockoff Filter (KO). In Figure 3, for each simulation we plot the power of each method as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$. The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of runs used in each simulation. Similarly, in Supplementary Figure S2, we plot the estimated FDR: the proportion of irrelevant variables in the discovery list, averaged over the runs.
Figure 3: RESET Ensemble vs. AdaKO GAM vs. KO. Each panel plots the power for each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction. A description of AdaKO GAM and KO can be found in Section 2.
@@
Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating that it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
@@
In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power. On the other hand, AdaKO GAM is more powerful than RESET at lower FDR thresholds, particularly in the case of $k=50$ and $\alpha=5\%$. This is not too surprising for two reasons. First, RESET trains its classification algorithms on a relatively small positive set of false nulls. Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overfits and incorrectly scores a knockoff higher than it should. Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff may cost up to $402/\alpha=$ 40 discoveries.
@@
We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO. The result in Simulation 3 is consistent with the findings of Ren and Candès in their analogous simulation. Indeed, AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not ‘easily’ separable. Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$. We hypothesize that the greater number of variables in this example, particularly relevant variables, appears to overcome the relatively poor performance of RESET Ensemble in Simulation 1 at low FDR thresholds.
Figure 4: Relative power and runtimes for each method. (A) Each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3. (B) The log10 runtimes of each method applied at the 5% FDR level for 5 runs of Simulation 4. For readability, the points are jittered in the horizontal direction.
We compared RESET Ensemble with AdaKO GAM because those are the respective default methods. We complement this comparison with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods. Specifically, for each panel and FDR threshold in Figure 3, we determined the maximum power across all methods and computed the relative power of each method to this maximum. The resulting relative powers are reported as boxplots, giving an indication of the performance of each method across the simulations as a whole. A plot of the actual powers for each of the methods is given in Supplementary Figure S3. While in Supplementary Figure S3 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions. On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated. Interestingly, RESET NN seems to be marginally more powerful than RESET Ensemble. While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any single classification algorithm.
One advantage of RESET is its computational speed. Indeed, our implementation of RESET, as described in Section 4.3, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores. In contrast, Adaptive Knockoffs must iteratively apply a classification algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed. Figure 4B demonstrates this advantage showing the (log10) runtimes of RESET (using 10 of a 20-cores CPU) and of Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4. We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs. RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-4 times faster than AdaKO LR and AdaKO GAM. In scenarios, as in Simulations 3 and 4, where AdaKO EM and AdaKO RF are infeasible, and AdaKO GAM and AdaKO LR are unable to leverage the side information RESET is clearly more suitable. We investigate such a commonly encountered scenario in the following application.
@@ 5.2 Application to peptide detection
In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR, and (c) when controlling the FDP, RESET is able to typically discover more peptides than the vanilla FDP-SD. We provide a heuristic justification of Assumption 3, which is required for RESET’s theoretical guarantees, in Section S.15.
5.2.1 HEK293 data
We show that the application of Adaptive Knockoffs is often too slow to be used effectively in scenarios such as the peptide detection problem that involve a large amount of data. We demonstrate this using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 50, 32] that we downloaded from the Proteomics Identification Database (PRIDE) [41]. After data preparation and searching the combined 24 HEK293 spectrum files, we determined the scores $Z_{i}$ for each target peptide and $\tilde{Z}_{i}$ for each decoy peptide using the XCorr score function (see Supplementary Section S.12 for details). Next, we competed each pair by recording the score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and label $L_{i}=1$ if $Z_{i}>\tilde{Z}_{i}$ and $L_{i}=-1$ if $Z_{i}<\tilde{Z}_{i}$ (ties were randomly broken). In addition to its primary score we associate with each peptide-spectrum match (PSM) a set of features, such as the charge state or peptide length. Each winning peptide inherits these features, or side information $\mathbf{x}_{i}$, from its maximally associated PSM, allowing us to define the desired triples $(L,W,\mathbf{x})$. A complete description of the side information used is given in Table S4 of the Supplement.

### v3

@@ 4.4 Semi-supervised approach
In this section we describe our specific framework for RESET’s Step (3), which can use any semi-supervised machine learning model. It consists of two main steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, presumably containing many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses. In Section C.2, we outline some additional heuristics that we employ to improve the procedure’s speed and power, as well as its applicability to dependent data.
  1. (i)
    We first define the negative and positive sets, as well as the features that are subsequently used by a collection of classification algorithms: the negative set is the set of all training decoys, the positive set is a subset of top scoring pseudo targets (Section C.2.3), and the features are the combined score and side information $\left(W,\mathbf{x}\right)$.
  2. (ii)
    We randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout). For each considered classification algorithm, and for each $k\in[K]$, we train the classification algorithm using the features of the hypotheses in the subsets of the positive and negative sets that are in the $K]\setminus\{k\}$ training folds, and then we apply the trained model to all the hypotheses in the $k$th test fold. We consider a random forest (RF), a generalized additive model (GAM), and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
  3. (iii)
    To reduce variability, we repeat Step (ii) $r$ times (we used $r=10$).
  4. (iv)
    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm A1, which is similar to SSS+ of Equation (1) but without the ‘+1’) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels. We set $c=c_{t}:=1-s\cdot(1-c_{0})$ (recall that $\mathbb{P}(L_{i}=1)\leq c_{0}$ under Assumption 3). This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo targets to training decoys (note, this is not a requirement of FDR/FDX control).
  5. (v)
@@
  6. (vi)
    Using the chosen classification algorithm, we assign new scores $\widetilde{W}$ to the hypotheses. $\widetilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis in the test fold that it appears in, taken over the $r$ repetitions executed in Step (iii).
  7. (vii)
[…]
  8. (viii)
    The positive set is redefined as those pseudo targets that are ‘discovered’ in Step (vii). Steps (ii)-(vi) are then repeated with the new positive set.
  9. (ix)
    The final scores $\widetilde{W}$ are then reported.
We refer to the general RESET wrapper in combination with the above semi-supervised approach as RESET Ensemble. We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step (ii) is reduced to a random forest, generalized additive model, and a collection of neural networks as described, respectively.
@@ 4.5 Code availability
Our code is available at https://github.com/freejstone/stat_RESET_paper_code with an Apache-2.0 licence. We intend on developing a package for RESET in R.
@@ 5 Experiments: competition setting
In the following sections, we explore two applications of competition-based multiple testing. The first is in selecting relevant variables that are conditionally associated with a response. The second is in the detection of peptides from mass spectrometry data.
5.1 Variable selection
We conducted a wide range of simulations in the variable selection problem, where the goal is to identify relevant variables from $\{X_{i}:i=1,\dots,p\}$ that are conditionally associated to the response $y$. The variable selection problem uses a competition framework called the knockoff filter (Barber and Candès (2015); Candès et al. (2018)). Briefly, each variable $X_{i}$ is paired with an artificial knockoff variable, denoted as $\tilde{X}_{i}$. Each variable and knockoff is scored according to how well it correlates with the response $y$, producing $Z_{i}$ and $\tilde{Z}_{i}$, respectively. Then following Section 2, the pairs of scores can be converted into a winning score $W_{i}$ and a label $L_{i}$, and a competition-based multiple testing procedure can be applied (see Section C.3 for details).
@@ 5.1 Variable selection
The following simulations, detailed in Section C.5, are based on those introduced by Ren and Candès (2023). Each setup, is designed to satisfy Assumption 3, and in each we compare RESET Ensemble with Adaptive Knockoffs (AdaKO, with its default GAM filter, see Section C.4 for details), as well as with the generic Knockoff Filter (KO) focusing on FDR control.
  • In simulation 1 we consider a linear model with $n=1000$ observations with $p=900$ variables. We use one dimensional side-information $\mathbf{x}_{i}$ where lower values correspond to a greater chance of being a relevant feature ($\beta_{i}\neq 0$). We vary the number of relevant variables $k\in\{50,150,300\}$ and generate and analyze 100 datasets for each.
  • In simulation 2 we consider a larger linear model with $n=2000$ and $p=1800$ with two-dimensional side information. The side information is drawn from a joint normal distribution, with different mean vectors for the relevant and irrelevant variables respectively. We vary the number of relevant variables $k\in\{150,300,450\}$ and generate and analyze 20 datasets for each.
  • In simulation 3 we consider a high-dimensional logistic model with $n=1000$ observations and $p=1600$ variables. Relevant and irrelevant variables are determined according to the spatially dependent two-dimensional side information $\mathbf{x}_{i}=(\mathbf{x}_{i1},\mathbf{x}_{i2})$ (see Figure C9). We generate and analyze 100 datasets.
  • In simulation 4 we consider a linear model with $n=10K$ observations and $p=10K$ variables with $k=0.15p$ relevant variables and three dimensional side information $\mathbf{x_{i}}=(\mathbf{x}_{i1},\mathbf{x}_{i2},\mathbf{x}_{i3})$. Lower values of the $\mathbf{x}_{ij}$’s correspond to a greater chance of being a relevant feature. We generate and analyze 5 datasets.
@@ 5.1.1 Results
In Figure 1, for each simulation we plot the power of each method as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$. The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of datasets drawn for each simulation. Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating that it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
Figure 1: Comparing the power of RESET Ensemble, AdaKO GAM and KO. Each panel shows the power of each method at FDR thresholds ranging from 5% to 30%. The first row corresponds to Simulation 1 with three values of $k\in\{50,150,300\}$, the second row corresponds to Simulation 2 with three values of $k\in\{150,300,450\}$, and the last row corresponds to Simulation 3 and 4. For readability, the points are jittered in the horizontal direction.
@@
In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power. On the other hand, AdaKO GAM is more powerful than RESET at lower FDR thresholds, particularly in the case of $k=50$ and $\alpha=5\%$. This is not too surprising for two reasons. First, RESET trains its classification algorithms on a relatively small positive set of false nulls. Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overfits and incorrectly scores a knockoff win higher than it should. Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff win may cost up to $2/\alpha=40$ discoveries.
@@
We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO. Indeed, The result in Simulation 3 is consistent with the findings of Ren and Candès in their analogous simulation. AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, effectively reducing to a linear model, and is therefore unable to leverage the side information if the true and false null hypotheses are not ‘easily’ separable. Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$. We hypothesize that the greater number of variables in this example, particularly relevant variables, appears to overcome the relatively poor performance of RESET Ensemble in Simulation 1 at low FDR thresholds.
We compared RESET Ensemble with AdaKO GAM because those are the respective default methods. In Section E.1, we complement this comparison where we summarize the performance of all RESET and AdaKO methods, and also analyze the runtimes of those methods.
@@ 5.2 Application to peptide detection
The end-to-end strategy for identifying peptides in an unknown sample can be broken into three steps: data acquisition, search, and competition-based multiple testing. We next provide a brief summary on each of these steps with details given in Section C.6.
During data acquisition, peptides are isolated and selected from a sample where each peptide generates a spectrum using a technique called liquid chromatography mass spectrometry (LC-MS/MS). In the search phase, each spectrum is searched against a database of peptides, called target peptides, for a possible candidate peptide that generated the spectrum. This search phase allows us to provide each target peptide with a score, $Z_{i}$, which quantifies evidence against the null hypothesis $H_{i}$ that ‘the target peptide is not in the sample’. Since not all target peptides are in the sample, each spectrum is searched against a database of artificially constructed peptides, called decoy peptides. The decoy peptides are obtained by randomly shuffling or reversing the target peptides. Each decoy peptide is therefore paired with a corresponding target peptide, and acquires a similar score from the search, $\tilde{Z}_{i}$.
In the final testing phase, we obtain a winning score, $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and a label $L_{i}$ indicating whether the $i$th target peptide ($L_{i}=1$) or its corresponding decoy peptide ($L_{i}=-1$) scored higher. We can then apply the SSS+ procedure — with $c=1/2$ and $\alpha$ typically set to the $1\%$ FDR threshold — to $L$ and $W$ to obtain a list of target peptides believed to be in the sample. Note that the SSS+ procedure (with $c=1/2$) is commonly referred to as target-decoy competition or TDC in the proteomics community.
In the following results, we show that (a) the Adaptive Knockoff methods are often computationally infeasible in this context, and (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR. In Section F.2, we provide a heuristic justification of Assumption 3, which is required for RESET’s theoretical guarantees. In terms of FDX control, RESET typically discovers more peptides than the side-information-oblivious FDP-SD (Section E.3). As outlined in Section 4.2, these comparative analyses of RESET Ensemble are new and were not undertaken in Freestone et al. (2025).
5.2.1 Runtime comparison
We argue that the application of Adaptive Knockoffs is often too slow to be used effectively in scenarios such as the peptide detection problem that involve a large amount of data. We demonstrate this using HEK293 (Chick et al. (2015)), a popular dataset used in benchmarking the development of new software (Kong et al. (2017); Yu et al. (2020); Lazear (2023)) that we downloaded from the Proteomics Identification Database (PRIDE) (Martens et al. (2005)). Details regarding the HEK293 dataset and the search phase analysis are given in Section D.1.
Table 2: The average runtimes for each method at 1% FDR on the HEK293 dataset.